OpenStudy (anonymous):

The sum of two consecutive odd integers is 100. Find the integers. 2n + 2 = 100 n + 4 = 100 2n + 1 = 100 2n + 3 = 100

OpenStudy (anonymous):

let n be the smaller of the two integers... the expression for the next consecutive integer is n+2....

OpenStudy (anonymous):

2n + 2 = 100 2n + 2 - 2 = 100 - 2 2n = 98 \[n=\frac{ 98 }{ 2 }\]n = 49

OpenStudy (anonymous):

2n+2=100?

Directrix (directrix):

@khrisangel Yes